OpenStudy (anonymous):

Which literal equations are equivalent to w = Fd? Choose all answers that are correct. (( I'll post answers in comments below. ))

OpenStudy (igreen):

I think they want you to solve for other variables? Like F and d?

OpenStudy (igreen):

Solve for \(F\): \(w = Fd\) Divide \(d\) to both sides: \(\dfrac{w}{d} = F\) Solve for \(d\): \(w = Fd\) Divide \(F\) to both sides: \(\dfrac{w}{F} = d\)

OpenStudy (anonymous):

A. d = fw B. \[d = \frac{ w }{ f }\] C . \[f = \frac{ w }{ d }\] D. \[f = \frac{ d }{ w }\]

OpenStudy (igreen):

B and C, they're identical to mine.

OpenStudy (anonymous):

mg ur awesome @iGreen

OpenStudy (igreen):

It's just like a regular equation, but you have variables instead of numbers.

OpenStudy (igreen):

Thanks. :)